OpenStudy (lilkg77):

HELP! A student found the slope of a line that passes through the points (2, 16) and (5,4) to be -1/4. What mistake did he make? A. He calculated the run over the rise. B. He forgot the negative sign. C. He used the reciprocal of the slope. D. He did not make a mistake. That is the correct slope.

OpenStudy (lilkg77):

Im positive the slope is correct but i wanna make sure

pooja195 (pooja195):

Well lets figure out the slope \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] \[\LARGE \frac{4-16=-12}{5-2=3}\] Now we need to simplify our fraction \[\huge~\rm~\frac{ -12/3=-4 }{ 3/3=1 }\] \[\huge~\rm~\frac{ -4 }{ 1 }=-4\]

OpenStudy (lilkg77):

Is -4 the same as -1/4? or does that not count

pooja195 (pooja195):

-4 is not the same as -1/4

OpenStudy (lilkg77):

So that means is answer was incorrect since the slope is -4 right?

pooja195 (pooja195):

Correct so where did he make his mistake?

OpenStudy (lilkg77):

Calculated the run over rise?

pooja195 (pooja195):

Correct :)

OpenStudy (lilkg77):

Thanks again! :)

pooja195 (pooja195):

np :)